# Intrinsic galaxy alignments in CAMELS simulations and the significant impact of baryon model

### v1

@@ 4.3 Dependence on simulation parameters
For each catalogue $i$, we compute the alignment amplitude $\langle w_{m+}^{(i)}\rangle_{r_{\perp}}$ and display these values in the six subplots in Figure 5, against simulation parameters. A logarithmic scaling is applied to $A_{\text{SN1}}$, $_{\text{SN2}}, AA$$_{\text{AGN1}}$, and $A_{\text{AGN2}}$ to ensure these parameters are uniformly distributed. We measure correlation with $\sigma_{8}^{2}$ because $w_{m+}$ is approximately proportional to $\sigma_{8}^{2}$, not $\sigma_{8}$, under the linear approximation described in Section 2 (Singh_2016). Linear fits are overlaid on scatter and density plots from the sample means to guide the eye.
@@
No significant correlation is detected between alignment amplitude $\langle w_{m+}^{(i)}\rangle_{r_{\perp}}$ and either $\log A_{\text{AGN1}}$ or $\log A_{\text{AGN2}}$. A lack of AGN dependence is likely due to the $(25\,h^{-1}\,\text{Mpc})^{3}$ simulation volume that limits the full range of large-scale modes that influence intrinsic alignments in cosmological surveys (Pandey2023). Other simulation studies (tenneti2015; Soussana2020) show that AGN feedback has negligible impact on matter-shape correlations but may impact galaxy-shape correlations especially when considering high-mass halos.
@@ 4.4 Parameter dependence of ratio $\frac{w_{m+}}{w_{mm}}$
We detect no significant correlation between measured $\langle\omega^{(i)}\rangle_{r_{\perp}}$ and $\sigma_{8}^{2}$ indicating that dividing by $w_{mm}$ removes the (linear) amplitude density fluctuation dependence on $w_{m+}$ in $\frac{w_{m+}}{w_{mm}}$. This result broadly follows predictions from NLA (see Appendix A).
@@ 4.6 Intrinsic ellipticity and alignment dependence
For parameter $\Omega_{m}$, Pearson values across all samples are more positive with $\langle\tilde{w}^{(i)}_{m+}\rangle_{r_{\perp}}$ than with $\langle w^{(i)}_{m+}\rangle_{r_{\perp}}$. We also identify a direct negative correlation between $\Omega_{m}$ and average ellipticity magnitude $\langle|\varepsilon|\rangle_{i}$. By removing the ellipticity magnitude, the negative $\tilde{w}^{(i)}_{m+}$ correlation with $\Omega_{m}$ is reduced. This suggests that the absolute ellipticity magnitude accounts for some of the $\Omega_{m}$ dependence in the correlation function $w^{(i)}_{m+}$. Previous studies (tenneti2015; Soussana2020) have shown that the scale dependence of alignment is unaffected by ellipticity magnitude. Our results are consistent with these findings: as shown in Figure 3, normalizing the ellipticity magnitude simply rescales the amplitude of the correlation function $w_{m+}(r_{\perp})$ by a constant factor.
@@ 5 Conclusions
In separating galaxy catalogues into two samples split by sSFR (at $10^{-10.5}$ /yr), we reveal that star forming galaxies (those above this cut) exhibit alignment amplitudes that are weaker by an order of magnitude compared to quiescent galaxies. This supports the established view that galaxy morphology and star formation activity modulate intrinsic alignments (hirata2007; mandelbaum2011; 2011A&A...527A..26J).
@@ Appendix A Cosmology dependence of the ratio under NLA
More generally, we can demonstrate the insensitivity of the theoretical ratio $\frac{\bar{w}_{m+}}{\bar{w}_{mm}}$ via numerical evaluation of Eq. 15. We use the CCL package (2019ApJS..242....2C) to calculate the theoretical spectra $P_{\delta I}(k)$ and $P_{\delta}(k)$ under the NLA model. The conclusions are not affected by choice of non-linear model (e.g. HALOFIT: 2003MNRAS.341.1311S; 2012MNRAS.420.2551B)
@@ Appendix B Density autocorrelation integral constraint correction
This form is standard  (e.g. Peebles:1980yev), but is still approximate at small scales. However, as we discuss below, our conclusions are not, in fact, sensitive to the exact value of $C$.

### v2

@@ 4.3 Dependence on simulation parameters
For each catalogue $i$, we compute the alignment amplitude $\langle w_{m+}^{(i)}\rangle_{r_{\perp}}$ and display these values in the six subplots in Figure 5, against simulation parameters. A logarithmic scaling is applied to $A_{\text{SN1}}$, $_{\text{SN2}}, AA$$_{\text{AGN1}}$, and $A_{\text{AGN2}}$ to ensure these parameters are uniformly distributed. We measure correlation with $\sigma_{8}^{2}$ because $w_{m+}$ is approximately proportional to $\sigma_{8}^{2}$, not $\sigma_{8}$, under the linear approximation described in Section 2 (Singh and Mandelbaum, 2016). Linear fits are overlaid on scatter and density plots from the sample means to guide the eye.
@@
No significant correlation is detected between alignment amplitude $\langle w_{m+}^{(i)}\rangle_{r_{\perp}}$ and either $\log A_{\text{AGN1}}$ or $\log A_{\text{AGN2}}$. A lack of AGN dependence is likely due to the $(25\,h^{-1}\,\text{Mpc})^{3}$ simulation volume that limits the full range of large-scale modes that influence intrinsic alignments in cosmological surveys (Pandey et al., 2023). Other simulation studies (Tenneti et al., 2015; Soussana et al., 2020) show that AGN feedback has negligible impact on matter-shape correlations but may impact galaxy-shape correlations especially when considering high-mass halos.
@@ 4.4 Parameter dependence of ratio $\frac{w_{m+}}{w_{mm}}$
We detect no significant correlation between measured $\langle\omega^{(i)}\rangle_{r_{\perp}}$ and $\sigma_{8}^{2}$ indicating that dividing by $w_{mm}$ removes the (linear) amplitude density fluctuation dependence on $w_{m+}$ in $\frac{w_{m+}}{w_{mm}}$. This result broadly follows predictions from NLA (see Appendix A) where most clustering dependence is removed by dividing by $w_{mm}$..
@@ 4.6 Intrinsic ellipticity and alignment dependence
For parameter $\Omega_{m}$, Pearson values across all samples are more positive with $\langle\tilde{w}^{(i)}_{m+}\rangle_{r_{\perp}}$ than with $\langle w^{(i)}_{m+}\rangle_{r_{\perp}}$. We also identify a direct negative correlation between $\Omega_{m}$ and average ellipticity magnitude $\langle|\varepsilon|\rangle_{i}$. By removing the ellipticity magnitude, the negative $\tilde{w}^{(i)}_{m+}$ correlation with $\Omega_{m}$ is reduced. This suggests that the absolute ellipticity magnitude accounts for some of the $\Omega_{m}$ dependence in the correlation function $w^{(i)}_{m+}$. Previous studies (Tenneti et al., 2015; Soussana et al., 2020) have shown that the scale dependence of alignment is unaffected by ellipticity magnitude. Our results are consistent with these findings: as shown in Figure 3, normalizing the ellipticity magnitude simply rescales the amplitude of the correlation function $w_{m+}(r_{\perp})$ by a constant factor.
@@ 5 Conclusions
In separating galaxy catalogues into two samples split by sSFR (at $10^{-10.5}$ /yr), we reveal that star forming galaxies (those above this cut) exhibit alignment amplitudes that are weaker by an order of magnitude compared to quiescent galaxies. This supports the established view that galaxy morphology and star formation activity modulate intrinsic alignments (Hirata et al., 2007; Mandelbaum et al., 2010b; Joachimi et al., 2011).
@@ Appendix A Cosmology dependence of the ratio under NLA
More generally, we can demonstrate the insensitivity of the theoretical ratio $\frac{\bar{w}_{m+}}{\bar{w}_{mm}}$ via numerical evaluation of Eq. 15. We use the CCL package (Chisari et al., 2019) to calculate the theoretical spectra $P_{\delta I}(k)$ and $P_{\delta}(k)$ under the NLA model. The conclusions are not affected by choice of non-linear model (e.g. HALOFIT: Smith et al., 2003; Bird et al., 2012)
@@ Appendix B Density autocorrelation integral constraint correction
This form is standard  (e.g. Peebles, 1980), but is still approximate at small scales. However, as we discuss below, our conclusions are not, in fact, sensitive to the exact value of $C$.